OpenStudy (anonymous):

Which is equivalent to (2x - 4) 3 = 18? A. 6x-12=18 B. 5x-4=18 C. 6x-4=18 D. 6x+12=18 First answer gets a medal ^^

OpenStudy (zzr0ck3r):

6x-12=18 is

OpenStudy (anonymous):

can you show how you go the answer if it's A?

OpenStudy (zzr0ck3r):

\((2x-4\))*3=2x*3-4*3=3*2*x-12=6x-12

OpenStudy (djcall11):